A car takes 2 hours to travel between two cities. If the average speed decreases 20km/hr it will take 30 minutes more. What is the distance between the two cities.

Appreciate your help.
With thanks
Ingrid

Answer by stanbon(75887)

A car takes 2 hours to travel between two cities. If the average speed decreases 20km/hr it will take 30 minutes more. What is the distance between the two cities.
--------------------------------------
Original DATA:
time = 2 hrs. ; distance = x km ; rate = d/t = x/2 kph
---------------------------------------------------------------
New DATA:
time = 2.5 hrs ; distance = x km ; rate = d/t == x/2.5 kph
---------------------------------------------------------------
Equations:
original speed - new speed = 20 kph
(x/2) - (x/2.5) = 20
(2.5x - 2x)/(2*2.5) = 20
0.5x = 100
x = 200 km (distance between the two cities.
===============================================
Cheers,
Stan H.